OpenStudy (anonymous):

Find the derivative of f(x) = 4 divided by x at x = 2.

OpenStudy (cggurumanjunath):

f(x)=(4/x) f"(x)=-4/(x^2) now put x=2 and tell me the answer.

OpenStudy (anonymous):

1

OpenStudy (cggurumanjunath):

no try again

OpenStudy (cggurumanjunath):

-4/(2^2)=-4/4=-1

OpenStudy (cggurumanjunath):

i hope it is clear ?

OpenStudy (anonymous):

How did it became negative?

OpenStudy (anonymous):

-4

OpenStudy (cggurumanjunath):

see derivative of 1/x=(-1/x^2) this is the property of differentiation.

OpenStudy (cggurumanjunath):